### v2

# CoTracker: It is Better to Track Together

Abstract
We introduce CoTracker, a transformer-based model that tracks dense points in a frame jointly across a video sequence. This differs from most existing state-of-the-art approaches that track points independently, ignoring their correlation. We show that joint tracking results in a significantly higher tracking accuracy and robustness.
We also provide several technical innovations, including the concept of virtual tracks, which allows CoTracker to track 70k points jointly and simultaneously. Furthermore, CoTracker operates causally on short windows (hence, it is suitable for online tasks), but is trained by unrolling the windows across longer video sequences, which enables and significantly improves long-term tracking.
We demonstrate qualitatively impressive tracking results, where points can be tracked for a long time even when they are occluded or leave the field of view. Quantitatively, CoTracker outperforms all recent trackers on standard benchmarks, often by a substantial margin. Code and model weights are available at: https://co-tracker.github.io/
Figure 1: CoTracker is a new dense point tracker. It can track 70k points jointly on a single GPU and exploit the correlation between different tracks to improve its performance. It has exceptional long-term tracking performance even in the presence of occlusions or when points leave the camera view (examples from DAVIS [32], the bottom row shows object tracks only).
1 Introduction
Establishing point correspondences is a fundamental problem in computer vision, necessary for many downstream tasks. Here, we are interested in estimating point correspondences in videos containing dynamic objects and a moving camera. Specifically, given one or more 2D points which are the projection of certain physical points of the underlying 3D scene, the goal is to find the location of the same physical points in all other frames of the video.
There are two variants of this problem. In optical flow, the objective is to estimate the velocity of all points within a video frame. This estimation is performed jointly for all points, but the motion is only predicted at an infinitesimal distance. In tracking, the goal is to estimate the motion of points over an extended period of time. For efficiency, tracking methods typically focus on a sparse set of points and treat them as statistically independent. Even recent techniques such as TAPIR [11] and PIPs++ [52], which employ modern architectures and can track points even in the presence of occlusions, model tracks independently. However, this approximation is crude as points are often strongly correlated (e.g., when they belong to the same object).
In this paper, we hypothesise that accounting for the correlation between tracked points can significantly improve performance. Thus, we introduce CoTracker, a new neural tracker that, compared to prior deep networks for optical flow and point tracking, supports joint estimation of a very large number of tracks simultaneously. CoTracker outperforms prior trackers on all common tracking benchmarks by substantial margins, is based on a simple and extensible architecture, and is as flexible as prior solutions for single-point tracking, meaning that it can track arbitrary points, selected at any spatial location and time in the video, and support arbitrary configurations of such points.
We show that tracking more points jointly significantly improves tracking metrics. In fact, we also show that injecting additional support points (i.e., points that are not requested by the user but help the tracker) can further improve performance, and we study several support configurations.
CoTracker achieves its high performance also due to several other innovations in its design and learning formulation. The network is a transformer operating in a sliding window fashion on a two-dimensional representation of tokens, where dimensions are time and the set of tracked points. Via suitable self-attention operators, the transformer can consider each track as a whole for the duration of a window and can exchange information between tracks, thus exploiting their correlation. When tracking a very large number of points, furthermore, CoTracker introduces in its transformer design a small number of representative virtual track tokens, switching from expensive self-attention between virtual tracks to cross-attention between the latter and the real ones. This way, CoTracker can jointly track a near-dense set of tracks on a single GPU.
Another important contribution is unrolled training. CoTracker is an online algorithm that processes videos in a serial fashion via a sliding window. Within a window, the tracks are initialized with queried points, and the network is tasked with progressively refining these initial estimations through iterative applications of the transformer. Each subsequent overlapping window starts with refined predictions of the previous window and updates tracks for the new frames. We mask out points that have not yet been initialized in the current sliding window, which allows initializing tracks from any point in the video, including in the middle of a window, and to join windows together. We show that, by training the network by unrolling several sliding windows, CoTracker achieves excellent long-term tracking performance as well.
We train CoTracker on the synthetic TAP-Vid-Kubric [10] and evaluate it on three different benchmarks, TAP-Vid-DAVIS, PointOdyssey [52] and DynamicReplica [23]. Our architecture works well for tracking single points and excels for groups of points, obtaining state-of-the-art tracking performance in several benchmarks.
2 Related work
Optical flow.
Optical flow estimates dense instantaneous motion. Originally approached by studying the colour constancy equation [17, 27, 3, 4], optical flow is nowadays tackled by using deep learning, starting with FlowNet [12, 19]. More recently, DCFlow [48] has introduced the computation of a 4D cost volume, used in most follow-up works [41, 45, 43]. Notable is RAFT [43], which introduced incremental flow updates and inspired a wave of subsequent works [22, 47, 22, 49]. We also adopt 4D cost volumes and iterated updates, but apply them to tracking.
Transformers [44] have also been applied to the optical flow problem [18, 36, 50]. Flowformer [18] drew inspiration from RAFT and proposed a transformer-based approach that tokenizes the 4D cost volume. GMFlow [50] replaced the update network with a softmax with self-attention for refinement. Perceiver IO [20] proposed a unified transformer for several tasks including optical flow.
Optical flow can in principle be applied to tracking by integrating predictions over time, but this open-loop approach accumulates errors in the form of drift, which motivates developing dedicated architectures like ours.
Multi-frame optical flow.
There have been several approaches [33, 35, 21, 28] to extend optical flow to multiple frames. Traditionally, Kalman filtering [9, 13] was a staple mechanism to ensure temporal consistency. Modern multi-frame methods produce dense flow. RAFT [43] can be applied in a warm-start manner [39, 42] for multi-frame optical flow estimation. However, these methods are not designed for long-term tracking and do not consider points occluded for a long time. VideoFlow [35] extends optical flow to three and five consecutive frames by integrating forward and backward motion features during flow refinement. MFT [31] conducts optical flow estimation between distant frames and chooses the most reliable chain of optical flows. Our tracker scales to produce semi-dense tracks.
Figure 2: CoTracker architecture. We compute convolutional features $\phi(I_{t})$ for every frame and process them with sliding windows. In order to initialize track features $Q$, we bilinearly sample from $\phi(I_{t})$ with starting point locations $P$. Locations $P$ also serve to initialize estimated tracks $\hat{P}$. See Fig. 3 for a detailed visualization of one sliding window.
Joint Tracking.
Before the emergence of deep learning, some authors proposed handcrafted joint trackers [2, 34], but there is less work on doing so with modern deep network architectures like ours. Our work is weakly related to multiple object tracking [8] where tracking through occlusions [51], with changes in appearance [29], and with temporal priors [37] have been extensively studied. However, our focus is the tracking of points, not objects.
Tracking any point.
TAP-Vid [10] introduced the problem of tracking any physical point in a video and proposed a benchmark and a simple baseline method for it, but is unable to track occluded points. PIPs [15] revisits the classic Particle Video [34] problem by introducing a model for point tracking through occlusions. The method tracks selected points with a fixed sliding window and restarts from the last frame where the point is visible. However, the model loses the target if it stays occluded beyond the size of the window. While the original Particle Video does track points jointly, PIPs and TAP-Vid track points in parallel, but independently. TAPIR [11] proposes a feed-forward point tracker with a matching stage inspired by TAP-Vid [10] and a refinement stage inspired by PIPs [15], which further improves tracking. PointOdyssey [52] addresses long-term tracking with PIPs++, a simplified version of PIPs, and introduces a benchmark for long-term tracking. However, both of these methods still track points independently. OmniMotion [46] optimizes a volumetric representation for each video, refining estimated correspondences in a canonical space. However, this approach requires test-time optimization, which, due to its cost, is not suitable for many practical applications, especially online ones.
Trackers and optical flow estimators are often trained using synthetic datasets [24, 30, 12, 40, 5, 52], as annotating real data can be challenging. Synthetic datasets provide accurate annotations, and training on them has demonstrated the ability to generalize to real-world data [30, 40].
3 CoTracker
Figure 3: CoTracker architecture. Visualization of one sliding window with $M$ iterative updates. During one iteration, we update point tracks $\hat{P}^{(m)}$ and track features $Q^{(m)}$. $Q^{(0)}$ is initialized with the initially sampled features $Q$ for all sliding windows, $\hat{P}^{(0)}$ with the starting locations for the first window. For other windows, $\hat{P}^{(0)}$ starts with predictions for frames processed in the preceding sliding window, and with the last predicted positions for the unseen frames. We compute visibility $\hat{v}$ after the last update $M$.
Our goal is to track 2D points throughout the duration of a video $V=(I_{t})_{t=1}^{T}$, which is a sequence of $T$ RGB frames $I_{t}\in\mathbb{R}^{3\times H\times W}$. The goal of the tracker is to predict $N$ point tracks $P^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $t=t^{i},\dots,T$, $i=1,\dots,N$, where $t^{i}\in\{1,\dots,T\}$ is the time when the track starts. The tracker also predicts the visibility flag $v^{i}_{t}\in\{0,1\}$ which tells if a point is visible or occluded in a given frame. To make the task definition unambiguous [10], we assume each point is visible at the start of their track (i.e., $v^{i}_{t_{i}}=1$). The tracker is thus given as input the video $V$ and the starting locations and times $(P^{i}_{t^{i}},t^{i})_{i=1}^{N}$ of $N$ tracks, and outputs an estimate $(\hat{P}^{i}_{t}=(\hat{x}^{i}_{t},\hat{y}^{i}_{t}),\hat{v}^{i}_{t})$ of the track locations and visibility for all valid ($t\geq t^{i}$) times.
3.1 Transformer formulation
We implement the tracker as a transformer neural network (Figs. 2 and 3) $\Psi:G\mapsto O$. The goal of this transformer is to improve an initial estimate of the tracks. Tracks are encoded as a grid of input tokens $G^{i}_{t}$, one for each track $i=1,\dots,N$, and time $t=1,\dots,T$. The updated tracks are expressed by a corresponding grid of output tokens $O^{i}_{t}$.
Image features.
We extract dense $d$-dimensional appearance features $\phi(I_{t})\in\mathbb{R}^{d\times\frac{H}{k}\times\frac{W}{k}}$ from each video frame $I_{t}$ using a convolutional neural network (trained end-to-end). We reduce the resolution by $k=4$ for efficiency. We also consider several scaled versions $\phi_{s}(I_{t})\in\mathbb{R}^{d\times\frac{H}{k2^{s-1}}\times\frac{W}{k2^{s-1}}}$, of the features with strides $s=1,\dots,S$ and use $S=4$ scales. These downscaled features are obtained by applying average pooling to the base features.
Track features.
The appearance of the tracks is captured by feature vectors $Q_{t}^{i}\in\mathbb{R}^{d}$ (these are time-dependent to accommodate changes in the track appearance). They are initialized by sampling image features at the starting locations, and then they are updated by the neural network, as explained below.
Correlation features.
In order to facilitate matching tracks to images, we adopt the correlation features $C^{i}_{t}\in\mathbb{R}^{S}$ of RAFT [43]. Each $C^{i}_{t}$ is obtained by comparing the track features $Q_{t}^{i}$ to the image features $\phi_{s}(I_{t})$ around the current estimate $\hat{P}_{t}^{i}$ of the tracks’s location. Specifically, the vector $C_{t}^{i}$ is obtained by stacking the inner products $[C_{t}^{i}]_{s\delta}=\langle Q^{i}_{t},~{}\phi_{s}(I_{t})[\hat{P}_{t}^{i}/ks+\delta]\rangle,$ where $s=1,\dots,S$ are the feature scales and $\delta\in\mathbb{Z}^{2}$, $\|\delta\|_{\infty}\leq\Delta$ are offsets. The image features $\phi_{s}(I_{t})$ are sampled at non-integer locations by using bilinear interpolation and border padding. The dimension of $C^{i}_{t}$ is $\big{(}2\Delta+1)^{2}S=196$ for our choice $S=4;\Delta=3$.
Tokens.
The input tokens $G(\hat{P},\hat{v},Q)$ code for position, visibility, appearance, and correlation of the tracks and is given by the following concatenation of features:
$G^{i}_{t}=(\hat{P}^{i}_{t}-\hat{P}^{i}_{1},~{}\hat{v}^{i}_{t},~{}Q^{i}_{t},~{}C^{i}_{t},~{}\eta(\hat{P}^{i}_{t}-\hat{P}^{i}_{1}))+\eta^{\prime}(\hat{P}^{i}_{1})+\eta^{\prime}(t).$ (1)
All the components except the last one have been introduced above. The last component is derived from the estimated position: it is the sinusoidal positional encoding $\eta$ of the track location with respect to the initial location at time $t=1$. We also add encodings $\eta^{\prime}$ of the start location $P^{i}_{1}$ and for the time $t$, with appropriate dimensions. In fact, we found it beneficial to separately encode the position of points at the first frame and their relative displacement to this frame.
The output tokens $O(\Delta\hat{P},\Delta Q)$ contain updates for location and appearance, i.e. $O^{i}_{t}=(\Delta\hat{P}^{i}_{t},\Delta Q^{i}_{t}).$
Iterated transformer applications.
We apply the transformer $M$ times in order to progressively improve the track estimates. Let $m=0,1,\dots,M$ index the estimate, with $m=0$ denoting initialization. Each update computes
$O(\Delta\hat{P},\Delta Q)=\Psi(G(\hat{P}^{(m)},\hat{v}^{(0)},Q^{(m)})).$
and sets $\hat{P}^{(m+1)}=\hat{P}^{(m)}+\Delta\hat{P}$ and $Q^{(m+1)}=Q^{(m)}+\Delta Q.$ The visibility mask $\hat{v}$ is not updated by the transformer, but once after the last of the $M$ applications of the transformer as $\hat{v}^{(M)}=\sigma(WQ^{(M)})$, where $\sigma$ is the sigmoid activation function and $W$ is a learned matrix of weights. We found iterative updates for the visibility did not further improve the performance, likely due to the fact that visibility highly depends on predicting an accurate location first.
For $m=0$, the estimates $\hat{P}^{(0)}$, $v^{(0)}$ and $Q^{(0)}$ are initialized by broadcasting the initial values $P^{i}_{t^{i}}$ (the location of query point), $1$ (meaning visible) and $\phi(I_{t^{i}})[P^{i}_{t^{i}}/k]$ (the appearance of the query point) to all times $t=1,\dots,T$.
Figure 4: Tracked and support points. We visualize different configurations of points along with their tracks. Points represent the start of each track, which clearly illustrates grids of points in (c) and (d). (a) A single target point in a sequence from TAP-Vid-DAVIS. (b) All target points — observe that they tend to cluster on objects of interest, potentially giving our tracker an ’unfair’ advantage compared to single-track methods. (c) A local grid of support points. (d) A global grid of support points. (e) SIFT support points. We evaluate CoTracker by combining different configurations of support points, as described in Sec. 3.5
3.2 Transformer architecture and virtual tracks
The transformer $\Psi$ interleaves attention operators that operate across the time and track dimensions, respectively. Factorising the attention [1] across time and tracks makes the model computationally tractable: the complexity is reduced from $O(N^{2}T^{2})$ to $O(N^{2}+T^{2})$. However, for very large values of $N$, this cost is still prohibitive, especially when tracking points densely in a video. We thus propose a new design, in which we introduce $K$ virtual tracks tokens, where $K\ll N$ is a hyperparameter. These are just added as additional tokens to the input grid with a fixed, learnable initialization and removed at the output. They participate in time attention as regular tracks, and in cross attention with all other tracks, reducing the cost to $O(NK+K^{2}+T^{2})$.
Starting with DETR [6], various computer vision transformers have used learnable queries. For instance, [20] uses them to decode dense outputs. Our virtual tracks are different, however: they mirror and shadow the computation of the prediction targets (i.e., the tracks), actively participating in the calculations to reduce the computational burden.
3.3 Windowed inference
A key advantage of formulating tracking as iterated updates is that it can easily support a sliding window application to process long videos. Consider in particular a video $V$ of length $T^{\prime}>T$ longer than the maximum window length $T$ supported by the architecture. To track points throughout the entire video $V$, we split the video in $J=\left\lceil 2T^{\prime}/T-1\right\rceil$ windows of length $T$, with an overlap of $T/2$ frames. [^1] We assume that $T$ is even. The last window is shorter if $T/2$ does not divide $T^{\prime}$.
Let the superscript $(m,j)$ denote the $m$-th application of the transformer to the $j$-th window. We thus have a $M\times J$ grid of quantities $(\hat{P}^{(m,j)},~{}\hat{v}^{(m,j)},~{}Q^{(m,j)})$, spanning transformer iterations and windows. Quantities $m=0$ and $j=1$ are initialized as for the single window case. Then, the transformer is applied $M$ times to obtain the estimate $(M,1)$. The latter is used to initialize estimate $(0,2)$ by broadcasting the known quantities as before. Specifically, the first $T/2$ components of $\hat{P}^{(0,2)}$ are copies of the last $T/2$ components of $\hat{P}^{(M,1)}$; the last $T/2$ components of $\hat{P}^{(0,2)}$ are instead copies of the last time $t=T/2-1$ from $\hat{P}^{(M,1)}$. The same update rule is used for $\hat{v}^{(0,2)}$, while $Q^{(0,j)}$ is always initialized with initial track features $Q$. This process is repeated until estimate $(M,J)$ is obtained.
3.4 Unrolled window training
We found it important to learn the windowed transformer in an unrolled fashion in order to learn to handle semi-overlapping windows. The primary loss is for track regression, summed over iterated transformer applications and windows:
$\mathcal{L}_{1}(\hat{P},P)=\sum_{j=1}^{J}\sum_{m=1}^{M}\gamma^{M-m}\|\hat{P}^{(m,j)}-P^{(j)}\|,$ (2)
where $\gamma=0.8$ discounts early transformer updates. Here $P^{(j)}$ contains the ground-truth trajectories restricted to window $j$ (trajectories which start in the middle of the window are padded backwards). The second loss is the cross entropy of the visibility flags $\mathcal{L}_{2}(\hat{v},v)=\sum_{j=1}^{J}\operatorname{CE}(\hat{v}^{(M,j)},v^{(j)})$. While only a moderate number of windows are used in the loss during training due to the computational cost, at test time we can unroll the windowed transformer applications arbitrarily, thus in principle handling any video length.
Implementation details.
A technical difficulty is that the number of tracks $N$ is not fixed, but varies from window to window as new points can be added to the tracker at any time. While conceptually this is handled by simply adding more tokens when needed, in practice changing the number of tokens makes batching data for training difficult. Instead, we allocate enough tokens to support all points that must be tracked, regardless of which window they are added in, and use masking to ignore tokens that are not yet used.
3.5 Support points
CoTracker can take advantage of tracking several points jointly. It is typical for applications to have several points to track, but in some cases, one might be interested in tracking a few or even a single point. In fact, single-point tracking is a case we consider in the evaluation for fairness (Sec. 1).
In practice, we found it beneficial to track additional ‘support points’ which are not explicitly requested by the user, particularly when the goal is to track only a few of them. Moreover, we have found that different configurations of such support points can lead to small differences in performance. We experiment with various configuration types, visualized in Fig. 4. With the “global” strategy, the support points form a regular grid across the whole image. With the “local” strategy, the grid of points is centred around the point we wish to track, thus allowing the model to focus on a neighbourhood of it. We also test using the SIFT detector [26] to select support points. Note that these patterns are only considered at inference time and are used to improve the accuracy of the tracker for the target points.
4 Experiments
We thoroughly evaluate CoTracker on standard real and synthetic tracking benchmarks to assess its performance and generalization properties in difficult conditions and against numerous state-of-the-art trackers.
Method Training data DAVIS First DAVIS Strided Dynamic Replica
AJ $\uparrow$ $\delta_{\text{avg}}^{\text{vis}}$uparrow OA $\uparrow$ AJ $\uparrow$ $\delta_{\text{avg}}^{\text{vis}}$uparrow OA $\uparrow$ $\delta_{\text{avg}}$uparrow $\delta_{\text{avg}}^{\text{vis}}$uparrow $\delta_{\text{avg}}^{\text{occ}}$uparrow
TAP-Net [10] TAP-Vid-Kubric 33.0 48.6 78.8 38.4 53.1 82.3 45.5 53.3 20.0
OmniMotion [46] 51.7 67.5 85.3
PIPs [15] FlyingThings++ 42.2 64.8 77.7 52.4 70.0 83.6 41.0 47.1 21.0
MFT [31] Kubric + others 47.3 66.8 77.8 56.1 70.8 86.9
PIPs++ [52] PointOdyssey 69.1 73.7 55.5 64.0 28.5
TAPIR [11] TAP-Vid-Kubric 56.2 70.0 86.5 61.3 73.6 88.8 56.8 66.1 27.2
CoTracker (Ours) TAP-Vid-Kubric 62.2 75.7 89.3 65.9 79.4 89.9 61.6 68.9 37.6
Table 1: State of the art comparison. We compare CoTracker to the best trackers available on TAP-Vid-DAVIS utilizing both their “queried first” and “queried strided” evaluation protocols as well as on Dynamic Replica. During evaluation on DAVIS, TAPIR and PIPs++ have access to all the video frames at once, while we process videos in an online manner, using a causal sliding window.
Method PointOdyssey
$\delta_{\text{avg}}$uparrow $\delta_{\text{avg}}^{\text{vis}}$uparrow $\delta_{\text{avg}}^{\text{occ}}$uparrow Survival $\uparrow$
TAP-Net [10] 28.4 18.3
PIPs  [16] 27.3 42.3
${}^{\dagger}$  [52] 29.0 32.4 18.8 47.0
CoTracker (Ours) 30.2 32.7 24.2 55.2
Table 2: PointOdyssey. All methods are trained on PointOdyssey. ${}^{\dagger}$ results for [52] obtained using released code and model.
Attention Points DAVIS First
time joint target
support
AJ $\uparrow$ $\delta_{\text{avg}}$uparrow OA $\uparrow$
(a) single
58.3 71.5 86.4
(b) single
41.1 62.9 75.1
(c) single
glob. $9\!\times\!9$
56.8 71.2 85.8
(d) single
SIFT $8\!\times\!8$
57.4 72.1 85.8
(e) single
loc. $10\!\times\!10$
60.4 75.2 87.5
(f) single
glob. $5\!\times\!5$ +loc. $8\!\times\!8$
62.2 75.7 89.3
(g) all
57.6 73.2 86.1
(h) all
glob. $5\!\times\!5$
60.5 75.8 87.6
(i) all
SIFT $8\!\times\!8$
60.7 75.7 88.1
Table 3: Importance of joint tracking. We compare using time and cross-track attention, tracking single or multiple target points, and using additional support points.
Figure 5: Qualitative Results. For PIPs++ (top), many points are incorrectly tracked and end up being ‘stuck’ on the front of the car. TAPIR (middle) works well for visible points but fails to handle occluded points. As soon as a point becomes occluded, it starts moving chaotically around the image. Our CoTracker (bottom) produces cleaner tracks. The tracks are also more ‘linear’ than those of PIPs++ or TAPIR, which is accurate as the primary motion is a homography (the observer does not translate).
Unrolled Training DAVIS First
AJ $\uparrow$ $\delta_{\text{avg}}$uparrow OA $\uparrow$
44.6 60.5 75.3
62.2 75.7 89.3
Table 4: Unrolled training. CoTracker is built for sliding window predictions. Using them during training is important.
Table 5: The virtual tracks allow CoTracker to scale. We report the maximum number of tracks that can fit on a 80 GB GPU.
Table 6: Single point vs. joint tracking. We qualitatively compare tracking individual points (top row) to tracking all points together (bottom). Background points are colored cyan while foreground points are magenta. We find that single-point tracking leads to background points following the object motion, as the model cannot reason well about the global context of the video (cyan points in later frames). Tracking points together results in better tracks for foreground and background points.
Num. virt. tracks DAVIS First Max. num. tracks Time [s]
[…]
0 61.6 75.6 88.3 $97\!\times\!97$ 207.3
32 60.2 74.4 88.5 $263\!\times\!263$ 26.8
64 62.2 75.7 89.3 $263\!\times\!263$ 27.9
128 60.9 74.8 88.4 $263\!\times\!263$ 30.1
Figure 6: Number of points. Accuracy for additional global grid and local grid points. We compute $\delta_{avg}^{vis}$ on TAP-Vid-DAVIS to find the optimal grid sizes. We use 25 & 64 because of speed and higher occlusion accuracy ($89.3$ vs. $88.7$).
Table 5: The virtual tracks allow CoTracker to scale. We report the maximum number of tracks that can fit on a 80 GB GPU.
Table 6: Single point vs. joint tracking. We qualitatively compare tracking individual points (top row) to tracking all points together (bottom). Background points are colored cyan while foreground points are magenta. We find that single-point tracking leads to background points following the object motion, as the model cannot reason well about the global context of the video (cyan points in later frames). Tracking points together results in better tracks for foreground and background points.